solve the differential equation dy/dx +(x/1-x^2)y=xy^1/2 , y(0)=1


1
Expert's answer
2021-04-13T09:40:15-0400

"y'+\dfrac{x}{1-x^2}y-x\sqrt{y}=0"

Let "y=u*v"

"u'v+v'u+\dfrac{x}{1-x^2}uv-x\sqrt{uv}=0"

"v(u'+\dfrac{x}{1-x^2}u)+v'u-x\sqrt{uv}=0"

Let first part equal to 0. Then:

"\begin{cases}
u'+\dfrac{x}{1-x^2}u=0\\
v'u-x\sqrt{uv}=0
\end{cases}"

"\begin{cases}
u'=-\dfrac{x}{1-x^2}u\\
v'u=x\sqrt{uv}
\end{cases}"

"\begin{cases}
\dfrac{du}{u}=-\dfrac{xdx}{1-x^2}\\
\dfrac{dv}{dx}u=x\sqrt{uv}
\end{cases}"

"\begin{cases}
\ln{u}=\ln{\dfrac{1}{\sqrt{1-x^2}}}\\
\dfrac{dv}{\sqrt{v}}=\dfrac{xdx}{\sqrt{u}}
\end{cases}"

"\begin{cases}
u=\dfrac{1}{\sqrt{1-x^2}}\\
\dfrac{dv}{\sqrt{v}}=x\sqrt[4]{1-x^2}dx
\end{cases}"

"\begin{cases}
u=\dfrac{1}{\sqrt{1-x^2}}\\
2\sqrt{v}=-\dfrac{2}{5}(1-x^2)^{5/4}+C
\end{cases}"

"\begin{cases}
u=\dfrac{1}{\sqrt{1-x^2}}\\
2\sqrt{v}=-\dfrac{2}{5}(1-x^2)^{5/4}+C
\end{cases}"

"\begin{cases}
u=\dfrac{1}{\sqrt{1-x^2}}\\
v=(\dfrac{1}{5}(1-x^2)^{5/4}-\dfrac{C}{2})^2
\end{cases}"

"y=\dfrac{1}{\sqrt{1-x^2}}*(\dfrac{1}{5}(1-x^2)^{5/4}-\dfrac{C}{2})^2"

"y(0)=1=>\\
1=\dfrac{1}{\sqrt{1-0^2}}*(\dfrac{1}{5}(1-0^2)^{5/4}-\dfrac{C}{2})^2"

"1=\dfrac{1}{\sqrt{1}}*(\dfrac{1}{5}(1)^{5/4}-\dfrac{C}{2})^2\\
1=(\dfrac{1}{5}-\dfrac{C}{2})^2=>\\
C=-\dfrac{8}{5} \space or\space С=\dfrac{12}{5}"

"y_1=\dfrac{1}{25\sqrt{1-x^2}}*((1-x^2)^{5/4}+4)^2\\
y_2=\dfrac{1}{25\sqrt{1-x^2}}*((1-x^2)^{5/4}-6)^2"